# Functional inequalities on the biased and restricted cube: an induction approach

### v2

@@ 4 Conclusion and open problems
The weak Harper’s edge-isoperimetric inequality (2) has several easy proofs based on induction. Once using conditional expectation to expand the total influence and by inductive assumption, it remains to check that the following two-point inequality for $a_{0},a_{1}\in(0,1)$
$a_{1}\log_{2}\left(\frac{1}{a_{1}}\right)+a_{0}\log_{2}\left(\frac{1}{a_{0}}\right)+|a_{1}-a_{0}|\geq(a_{1}+a_{0})\log_{2}\left(\frac{2}{a_{1}+a_{0}}\right),$
which is not hard to verify, using, e.g., concavity of the entropy function $H(x)=-x\log x-(1-x)\log(1-x)$ on $[0,1]$. It may be more amazing that we can prove the ‘sharp’ Harper’s edge-isoperimetric inequality by using entropy method, see Section 4.4 in [7]. Since numerous classical results in Boolean function analysis–including the hypercontractivity inequality–can be established via entropy methods, see for example [32], we are very interested in whether other methods, especially analytical methods (or spectral graph theory), can be used to prove tight Harper’s edge-isoperimetric inequality.
A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([10, Question 1.1]). As observed in [17], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$. Following the general relations between log-Sobolev and mixing time bound (e.g.,[6, 9]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
Our inductive two-point machinery is well suited for quadratic (L2) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly. However, in log-Sobolev inequalities, the relative entropy is a nonlinear, non-quadratic functional—this introduces challenges when applying the induction method to control it. Hence, to establish log-Sobolev inequalities in non-product spaces, further developing the induction method or looking for new methods is needed.

### v3

@@ 6 Conclusion and open problems
A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([DM2014, Question 1.1]). As observed in [fei2025spectral], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$. Following the general relations between log-Sobolev and mixing time bound (e.g.,[BT2006logsobolev, DS1996logsobolev]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
Our inductive two-point machinery is well suited for quadratic ($L^{2}$) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly. However, in log-Sobolev inequalities, the relative entropy is a nonlinear, non-quadratic functional—this introduces challenges when applying the induction method to control it. Hence, to establish log-Sobolev inequalities in non-product spaces, further developing the induction method or looking for new methods is needed.